Refer to the following financial statements and answer the questions:

Answer the following questions:

  1. Determine which items varied in constant proportion to sales between 20x6 and 20x7.
  2. Determine the rate of growth in sales that was achieved from year 20x6 to 20x7.
  3. What was the firm’s return on equity for 20x7? Can you calculate for 20x6?
  4. What was the firm’s external (additional) funding requirement determined to be for 20x? How was the funding obtained? (hint: compare the changes in debts and equities).
Income statement
20x6 20x7
Sales $1,200,000 $1,500,000
Cost of goods sold $750,000 $937,500
Gross margin $450,000 $562,500
Operating expense
Advertising expense $50,000 $62,500
Rent expense $72,000 $90,000
Sales commission $48,000 $60,000
Utilities $15,000 $18,750
EBIT $265,000 $331,250
Interest expense $106,000 $113,000
Taxable income $159,000 $218,250
Tax (35%) $55,650 $76,388
Net Income $103,350 $141,863
Dividend (40% payout) $41,340 $56,745
Change in return earnings $62,010 $85,118
Balance sheet
20x6 20x7
Assets
Cash $300,000 $375,000
Receivables $200,000 $250,000
Inventory $700,000 $875,000
Property, plant and equipment $1,800,000 $2,250,000
Total assets $3,000,000 $3,750,000
Liabilities
Payables $300,000 $375,000
Short-term debt (10 % interest) $500,000 $989,882
Long-term debt (7% interest) $800,000 $900,000
Shareholders' equity
Common stock $1,100,000 $1,100,000
Retained earnings $300,000 $385,118
Total liabilities and equity $3,000,000 $3,750,000
Answer #1

a)variation in sales from 20x6 to 20x7= 1500000-1200000= 300000

percentage change= 300000/1200000 = 25%

COGS increased by 25% from 750000 to 937500

gross margin increased by 25%

advertising expense, rent expense, sales commission and utilities expense also increased by 25%

ebit also increased by 25%

b) rate of growth of sales= 1500000-1200000= 300000

percentage growth= 300000/1200000*100= 25%

c) return on equity= net income/ shareholders equity

for 20x6= 103350/1100000

= 0.093 or 9.3%

for 20x7= 141863/1100000

=0.128 or 12.8%

d) firms additional funding sources in20x7

payables-75000

short term debt-489882

long term debt- 100000